# Optimal Tests of the Composite Null Hypothesis Arising in Mediation Analysis

### v2

@@ 2 PRELIMINARIES
Because we are taking the intersection of rejection regions of two size-$\alpha$ tests, the resulting test must necessarily be conservative. In fact, when $(\delta_{x}^{*},\delta_{y}^{*})=(0,0)$, the probability of being in the joint significance test rejection region is $\alpha^{2}$, well below the desired level $\alpha$. By the continuity of the rejection probability in $(\delta_{x}^{*},\delta_{y}^{*})$, the rejection probability is likewise well below $\alpha$ in a region around the origin, hence the underpoweredness of the joint significance test. That is, the worst-case type 2 error of the joint significance test is
$\sup_{(\delta^{*}_{x},\delta^{*}_{y}):\delta^{*}_{x}\delta^{*}_{y}\neq 0}\mathrm{Pr}_{\delta^{*}_{x},\delta^{*}_{y}}\{(Z^{x}_{*},Z^{y}_{*})\notin R\}=1-\min_{(\delta^{*}_{x},\delta^{*}_{y}):\delta^{*}_{x}\delta^{*}_{y}=0}\mathrm{Pr}_{\delta^{*}_{x},\delta^{*}_{y}}\{(Z^{x}_{*},Z^{y}_{*})\in R\}=1-\alpha^{2}.$
The rejection region corresponding to the delta method-based test has been shown to be contained in that of the joint significance test (Huang,, 2019; van Garderen and van Giersbergen, 2022a, ), and therefore the delta method-based test is uniformly more conservative than the joint significance test, hence it suffers from the same problem.
Our goal in the following two sections is to define a new test by constructing a rejection region that preserves type 1 error uniformly over the null hypothesis space, but improves power in the alternative hypothesis space, particularly in the region around the origin of the parameter space.
@@ 3 MINIMAX OPTIMAL TEST
In this section, we seek a test defined by a rejection region $R^{*}\subset\mathbb{R}^{2}$ satisfying $\sup_{(\delta^{*}_{x},\delta^{*}_{y}):\delta^{*}_{x}\delta^{*}_{y}=0}\allowbreak\mathrm{Pr}_{\delta^{*}_{x},\delta^{*}_{y}}\allowbreak\{(Z_{*}^{x},\allowbreak Z_{*}^{y})\allowbreak\in\allowbreak R^{*}\}\allowbreak=\allowbreak\alpha$ and the minimax optimality criterion
$\inf_{(\delta^{*}_{x},\delta^{*}_{y}):\delta^{*}_{x}\delta^{*}_{y}\neq 0}\mathrm{Pr}_{\delta^{*}_{x},\delta^{*}_{y}}\{(Z_{*}^{x},Z_{*}^{y})\in R^{*}\}=\sup_{R\subset\mathbb{R}^{2}}\inf_{(\delta^{*}_{x},\delta^{*}_{y}):\delta^{*}_{x}\delta^{*}_{y}\neq 0}\mathrm{Pr}_{\delta^{*}_{x},\delta^{*}_{y}}\{(Z^{x}_{*},Z^{y}_{*})\in R\}.$ (5)
In words, the rejection region $R^{*}$ generates a test with type 1 error $\alpha$ that achieves the minimax risk of the 0-1 loss function, i.e., that yields the largest worst-case power. By continuity of the power function in $(\delta^{*}_{x},\delta^{*}_{y})$, the minimax optimal power is upper bounded by $\alpha$, since the rejection probability can be at most $\alpha$ at any given value of $(\delta^{*}_{x},\delta^{*}_{y})$ in the null hypothesis space, and can be arbitrarily close to this rejection probability in a small enough neighborhood around that point, which will contain elements of the alternative hypothesis space. Moreover, if (5) is attainable, a minimax optimal test generated by the rejection region $R$ must satisfy
$\mathrm{Pr}_{\delta^{*}_{x},\delta^{*}_{y}}\{(Z_{*}^{x},Z_{*}^{y})\in R\}=\alpha\quad\text{for all }(\delta^{*}_{x},\delta^{*}_{y})\in\mathbb{R}^{2}\text{ such that }\delta^{*}_{x}\delta^{*}_{y}=0.$ (6)
A test satisfying (6) is known as a similar test (Lehmann et al.,, 2005). Similarity is an important property that is closely tied to uniformly most powerful unbiased tests in the classical hypothesis testing literature (e.g., see Chapter 4 of Lehmann et al., (2005)). In our case, it is important because a non-similar test must have type 1 error either greater than or less than $\alpha$ somewhere in the null hypothesis space. In the former case, the test will fail to preserve type 1 error. In the latter, by continuity of the rejection probability as a function of the true parameter, the rejection probability will necessarily be less than $\alpha$ in some neighborhood that also contains the alternative hypothesis space. Thus, a non-similar test will necessarily be a biased test and underpowered in a region of the alternative hypothesis space.
@@ 3.1 A Minimax Optimal and Similar Test for Unit Fraction $\alpha$
We now define a rejection region that generates a similar test for any unit fraction $\alpha$, i.e., any $\alpha$ such that $\alpha^{-1}$ is a positive integer, hence a test that satisfies the minimax optimality criterion (5). Let $a_{k}=\Phi^{-1}(k\alpha/2)$ for $k=0,\ldots,2/\alpha$, where $\Phi$ is the standard normal distribution function. Define $R_{mm}:=\{\bigcup_{k=1}^{2/\alpha}(a_{k-1},a_{k})\times(a_{k-1},a_{k})\}\cup\{\bigcup_{k=1}^{2/\alpha}(a_{k-1},a_{k})\times(-a_{k},-a_{k-1})\}.$ The region $R_{mm}$ for $\alpha=0.05$ is depicted in the plot in Figure 1(b). Clearly, $R_{mm}$ contains the rejection region of the joint significance test, hence it yields a test that is uniformly more powerful. A plot of its power function surface over $\mathbb{R}^{2}$ is available in the Supplementary Materials.
@@
All proofs are presented in the appendix. The essence of the proof of Theorem 1 is that when $\delta^{*}_{y}=0$, whatever level $Z^{x}_{*}$ takes, the conditional probability of $Z^{y}_{*}$ being in $R_{mm}$ given $Z^{x}_{*}$ equals $\alpha$, since the probability of being in each vertical interval is $\alpha/2$ for a standard normal distribution. The same holds for $Z^{x}_{*}$ when $\delta^{*}_{x}=0$ by symmetry.
@@
The rejection region $R_{mm}$ satisfies the minimax optimality criterion (5) with $\inf_{(\delta^{*}_{x},\delta^{*}_{y}):\delta^{*}_{x}\delta^{*}_{y}\neq 0}\mathrm{Pr}_{\delta^{*}_{x},\delta^{*}_{y}}\{(Z_{*}^{x},Z_{*}^{y})\in R_{mm}\}=\alpha$.
van Garderen and van Giersbergen, (2020) give a result stating that there is no similar test of $H_{0}$ against $H_{1}$ within a particular class of tests defined by a rejection region generated by rotating and reflecting a monotonically increasing function from zero to infinity. While Corollary 1 may appear to contradict this result, $R_{mm}$ does not in fact belong to the class considered by van Garderen and van Giersbergen, (2020), but rather belongs to a class that relaxes the restriction that the function must be monotonically increasing to allow for it to be monotonically nondecreasing (though the more recent version (van Garderen and van Giersbergen, 2022a, ) recognizes the existence of this function). In fact, the following theorem states that the test generated by $R_{mm}$ is the only similar test in this relaxed class up to a set of measure zero.
@@
Let $\mathcal{F}$ be the class of all monotonically nondecreasing functions mapping all $x$ from the nonnegative real numbers to $[0,x]$, and for any $f\in\mathcal{F}$, let $R_{f}$ be the region generated by taking all possible negations and permutations of the region $\{(x,y):y\in(f(x),x]\}$, or equivalently reflecting this region about the $x$, $y$, $y=x$, and $y=-x$ axes, and rotating it $\pi/2$, $\pi$, and $23\pi/$ 2 radians about the origin. The function $f_{mm}(x):=\sum_{k=1/\alpha}^{2/\alpha}a_{k-1}I(a_{k-1}\leq x<a_{k})$, which generates $R_{mm}$, is the unique function in $\mathcal{F}$ (up to a set of measure zero) that generates a similar test.
In addition to minimax optimality, this test has some other desirable properties. It has a simple, exact closed form, making it very fast and straightforward to implement. Additionally, it is nonrandom, and is symmetric with respect to negation and permutations. One less desirable property is that the test will reject for $(Z^{x}_{n},Z^{y}_{n})$ arbitrarily close to $(0,0)$, which we discuss further in Section 6. However, from Theorem 2 we infer that at least within the class of tests generated by $\mathcal{F}$, this is necessary in order to attain the best worst-case power. In fact, more generally, we have the following result.
@@
There is no similar test of $H_{0}$ against $H_{1}$ with a rejection region that is bounded away from $\{(\delta_{x}^{*},\delta_{y}^{*}):\delta_{x}^{*}\delta_{y}^{*}=0\}$.
Therefore, one cannot have a hypothesis test of $H_{0}$ that is both similar/unbiased and that fails to reject for all values of $(Z_{n}^{x},Z_{n}^{y})$ that are arbitrarily close to the null hypothesis space; there is a fundamental trade-off between these two properties.
@@ 3.2 Dealing With Non-Unit Fraction Values of $\alpha$
The rejection region $R_{mm}$ is only defined above for unit fractions $\alpha$. For other values of $\alpha$, the class of tests generated by $\mathcal{F}$ will not contain a similar test, though nonrandom similar tests may exist outside of this class. Tests at other values of $\alpha$ are typically of interest for two reasons: using multiple testing adjustment procedures and defining $p$-values. We discuss the latter in the Supplementary Materials. The restriction that $\alpha$ must be a unit fraction is not a limitation for the Bonferroni correction procedure, which involves dividing the familywise error rate $\alpha$ by the number of tests, provided the familywise error rate $\alpha$ itself is a unit fraction. However, it is a limitation for the Benjamini–Hochberg procedure, which involves multiplying the false discovery rate $\alpha$ by a sequence of rational numbers, only some of which are unit fractions. We discuss the Benjamini–Hochberg procedure in Section 5.
We now define an extension of the minimax optimal test to possibly non-unit fraction values of $\alpha\in(0,1)$. If $\alpha$ is not a unit fraction, then this extension will not yield a minimax optimal test or a similar test, as similar tests do not exist in this class of tests for such values. However, they do preserve type 1 error, and have type 1 error that approximates $\alpha$ under the least-favorable distribution in the null hypothesis very closely as $\alpha$ goes to zero. Let $b_{-1}:=0$ and $b_{k}:=\Phi^{-1}\{1-(\lfloor\alpha^{-1}\rfloor-k)\alpha/2\}$ for $k=0,\ldots,\lfloor\alpha^{-1}\rfloor$ such that $0=b_{-1}\leq b_{0}<b_{1}<\ldots<b_{\lfloor\alpha^{-1}\rfloor}=+\infty$. We define the rejection region $\alpha$ for any $\alpha\in(0,1)$ to be $R_{mm}^{\prime}:=\{\bigcup_{k=0}^{\lfloor\alpha^{-1}\rfloor}(b_{k-1},b_{k})\times(b_{k-1},b_{k})\}\cup\{\bigcup_{k=0}^{\lfloor\alpha^{-1}\rfloor}(b_{k-1},b_{k})\times(-b_{k},-b_{k-1})\}\cup\{\bigcup_{k=0}^{\lfloor\alpha^{-1}\rfloor}(-b_{k},-b_{k-1})\times(b_{k-1},b_{k})\}\cup\{\bigcup_{k=0}^{\lfloor\alpha^{-1}\rfloor}(-b_{k},-b_{k-1})\times(-b_{k},-b_{k-1})\}$. When $\alpha$ is a unit fraction, $R^{\prime}_{mm}$ coincides with $R_{mm}$, hence $R^{\prime}_{mm}$ generates the minimax optimal test discussed earlier. Otherwise, the test generated by $R^{\prime}_{mm}$ corresponds closely to the minimax optimal test in that if $\delta_{y}^{*}=0$, then given $Z^{x}_{*}=z$ with $z\notin[-b_{0},b_{0}]$, the probability of being in the rejection region is exactly $\alpha$. Precisely:
@@
$\lfloor\alpha^{-1}\rfloor\alpha^{2}+(1-\lfloor\alpha^{-1}\rfloor\alpha)^{2}$ $=\inf_{(\delta^{*}_{x},\delta^{*}_{y}):\delta^{*}_{x}\delta^{*}_{y}=0}\mathrm{Pr}_{\delta^{*}_{x},\delta^{*}_{y}}\{(Z_{*}^{x},Z_{*}^{y})\in R_{mm}^{\prime}\}$ (7)
$\leq\sup_{(\delta^{*}_{x},\delta^{*}_{y}):\delta^{*}_{x}\delta^{*}_{y}=0}\mathrm{Pr}_{\delta^{*}_{x},\delta^{*}_{y}}\{(Z_{*}^{x},Z_{*}^{y})\in R_{mm}^{\prime}\}=\alpha.$
@@ 3.3 Asymptotic approximation considerations
The results we have given so far have been in terms of the limiting distribution $(Z^{x}_{*},Z^{y}_{*})$. However, if the joint convergence of $(\hat{\delta}_{x},\hat{\delta}_{y})$ is uniform in $(\delta_{x},\delta_{y})$ as in (1), then the power function $\beta_{n}(\delta_{n}):=\mathrm{Pr}_{\hat{\delta}_{x},\hat{\delta}_{y}}\{(Z^{x}_{n},Z^{y}_{n})\in R_{mm}\}$ for the observed data distribution will converge uniformly to $\beta_{*}(\delta^{*}):=\mathrm{Pr}_{\delta_{x}^{*},\delta_{y}^{*}}\{(Z^{x}_{*},Z^{y}_{*})\in R_{mm}\}$, i.e., that of the limiting distribution. Thus, for any arbitrarily small tolerance $\epsilon>0$, there is an $N$ such that $\beta_{n}(\delta_{n})$ will be within $\epsilon$ of $\beta_{*}(\delta^{*})$ for all $(\delta_{x},\delta_{y})$ and $n\geq N$. In particular, this means violations of type 1 error can be made arbitrarily small uniformly over $H_{0}$ given a sufficiently large sample size. The asymptotic approximation can be improved for a given finite sample $n$ by replacing all instances of the normal distribution with a $t$ distribution with $(n-1)$ degrees of freedom.

### v3

@@ 2 PRELIMINARIES
Because we take the intersection of rejection regions of two size-$\alpha$ tests, the resulting test is necessarily conservative. In fact, when $\bm{\delta}^{*}=(0,0)^{\top}$, the probability of being in the joint significance test rejection region is $\alpha^{2}$. By continuity of the rejection probability in $\bm{\delta}^{*}$, the rejection probability is likewise well below $\alpha$ in a region around the origin, hence the underpoweredness of the joint significance test. That is, the worst-case type 2 error of the joint significance test is
$\sup_{\bm{\delta}^{*}:\delta^{*}_{x}\delta^{*}_{y}\neq 0}\mathrm{Pr}_{\bm{\delta}^{*}}\{(Z^{x}_{*},Z^{y}_{*})\notin R\}=1-\min_{\bm{\delta}^{*}:\delta^{*}_{x}\delta^{*}_{y}=0}\mathrm{Pr}_{\bm{\delta}^{*}}\{(Z^{x}_{*},Z^{y}_{*})\in R\}=1-\alpha^{2}.$
Huang, (2019) and van Garderen and van Giersbergen, (2024) showed that the delta method-based test’s rejection region is contained in that of the joint significance test, and therefore the former is uniformly more conservative than the latter. In the following two sections, we define new tests by constructing a rejection region that preserves type 1 error uniformly over the null hypothesis space, but improves power in the alternative hypothesis space, particularly near the origin of the parameter space.
@@ 3 MINIMAX OPTIMAL TEST
In this section, we seek a test defined by a rejection region $R^{*}\subset\mathbb{R}^{2}$ satisfying $\sup_{\bm{\delta}^{*}:\delta^{*}_{x}\delta^{*}_{y}=0}\allowbreak\mathrm{Pr}_{\bm{\delta}^{*}}\allowbreak\{(Z_{*}^{x},\allowbreak Z_{*}^{y})\allowbreak\in\allowbreak R^{*}\}\allowbreak=\allowbreak\alpha$ and the minimax optimality criterion
$\inf_{\bm{\delta}^{*}:\delta^{*}_{x}\delta^{*}_{y}\neq 0}\mathrm{Pr}_{\bm{\delta}^{*}}\{(Z_{*}^{x},Z_{*}^{y})\in R^{*}\}=\sup_{R\subset\mathbb{R}^{2}}\inf_{\bm{\delta}^{*}:\delta^{*}_{x}\delta^{*}_{y}\neq 0}\mathrm{Pr}_{\bm{\delta}^{*}}\{(Z^{x}_{*},Z^{y}_{*})\in R\}.$ (5)
The rejection region $R^{*}$ generates a test with type 1 error $\alpha$ achieving the minimax risk of the 0-1 loss function, i.e., yielding the largest worst-case power. By continuity of the power function in $\bm{\delta}^{*}$, the minimax optimal power is upper bounded by $\alpha$. Moreover, if (5) is attainable, a minimax optimal test generated by the rejection region $R$ must satisfy
$\mathrm{Pr}_{\bm{\delta}^{*}}\{(Z_{*}^{x},Z_{*}^{y})\in R\}=\alpha\quad\text{for all }\bm{\delta}^{*}\in\mathbb{R}^{2}\text{ such that }\delta^{*}_{x}\delta^{*}_{y}=0.$ (6)
A test satisfying (6) is known as a similar test. Similarity is an important property that is closely tied to uniformly most powerful unbiased tests in the classical hypothesis testing literature. In our case, it is important because a non-similar test must have type 1 error either greater than or less than $\alpha$ somewhere in the null hypothesis space. In the former case, the test will fail to preserve type 1 error. In the latter, the rejection probability will necessarily be less than $\alpha$ in some neighborhood that also contains the alternative hypothesis space. Thus, a non-similar test will necessarily be a biased test and underpowered in a region of the alternative hypothesis space.
@@ 3.1 A Minimax Optimal and Similar Test for Unit Fraction $\alpha$
We now define a rejection region that generates a similar test for any unit fraction $\alpha$, i.e., any $\alpha$ such that $\alpha^{-1}$ is a positive integer, hence a test that satisfies the minimax optimality criterion (5). Let $a_{k}=\Phi^{-1}(k\alpha/2)$ for $k=0,\ldots,2/\alpha$, where $\Phi$ is the standard normal distribution function. Define $R_{mm}:=\{\bigcup_{k=1}^{2/\alpha}(a_{k-1},a_{k})\times(a_{k-1},a_{k})\}\cup\{\bigcup_{k=1}^{2/\alpha}(a_{k-1},a_{k})\times(-a_{k},-a_{k-1})\}.$ The region $R_{mm}$ for $\alpha=0.05$ is depicted in the plot in Figure 1(b). Clearly, $R_{mm}$ contains the rejection region of the joint significance test, hence it yields a test that is uniformly more powerful.
@@
The essence of the proof of Theorem 1 is that when $\delta^{*}_{y}=0$ (resp. $\delta^{*}_{x}=0$), whatever level $Z^{x}_{*}$ (resp. $Z^{y}_{*}$) takes, the conditional probability of $Z^{y}_{*}$ (resp. $Z^{x}_{*}$) being in $R_{mm}$ given $Z^{x}_{*}$ (resp. $Z^{y}_{*}$) equals $\alpha$, since the probability of being in each vertical interval is $\alpha/2$ under the standard normal law.
@@
The rejection region $R_{mm}$ satisfies the minimax optimality criterion (5) with $\inf_{\bm{\delta}^{*}:\delta^{*}_{x}\delta^{*}_{y}\neq 0}\mathrm{Pr}_{\bm{\delta}^{*}}\{(Z_{*}^{x},Z_{*}^{y})\in R_{mm}\}=\alpha$.
van Garderen and van Giersbergen, (2020) gave a result stating that there is no similar test of $H_{0}$ against $H_{1}$ within a particular class of tests defined by a rejection region generated by rotating and reflecting a monotonically increasing function from zero to infinity. While Corollary 1 may appear to contradict this result, $R_{mm}$ does not in fact belong to the class considered by van Garderen and van Giersbergen, (2020), but rather belongs to a class that relaxes the restriction that the function must be monotonically increasing to allow for it to be monotonically nondecreasing. In fact, the following theorem states that the test generated by $R_{mm}$ is the only similar test in this relaxed class up to a set of measure zero.
@@
Let $\mathcal{F}$ be the class of all monotonically nondecreasing functions mapping all $x$ from the nonnegative real numbers to $[0,x]$, and for any $f\in\mathcal{F}$, let $R_{f}$ be the region generated by taking all possible negations and permutations of the region $\{(x,y):y\in(f(x),x]\}$, or equivalently reflecting this region about the $x$, $y$, $y=x$, and $y=-x$ axes, and rotating it $\pi/2$, $\pi$, and $23\pi/$ 2 radians about the origin. The function $x\mapsto f_{mm}(x):=\sum_{k=1/\alpha}^{2/\alpha}a_{k-1}I(a_{k-1}\leq x<a_{k})$, which generates $R_{mm}$, is the unique function in $\mathcal{F}$ (up to a set of measure zero) that generates a similar test.
This test has a simple, exact closed form, making it very fast and straightforward to implement. Additionally, it is nonrandom and symmetric with respect to negation and permutations. One less desirable property is that the test will reject for $(Z^{x}_{n},Z^{y}_{n})$ arbitrarily close to $(0,0)$, which we discuss further in Section 6. However, the following result shows this is necessary in order to attain the best worst-case power.
@@
There is no similar test of $H_{0}$ against $H_{1}$ with a rejection region that is bounded away from $\{\bm{\delta}^{*}:\delta_{x}^{*}\delta_{y}^{*}=0\}$.
Thus, one cannot have a hypothesis test of $H_{0}$ that is both similar/unbiased and that fails to reject for all values of $(Z_{n}^{x},Z_{n}^{y})$ that are arbitrarily close to the null hypothesis space.
@@ 3.2 Dealing With Non-Unit Fraction Values of $\alpha$
The rejection region $R_{mm}$ is only defined above for unit fractions $\alpha$. For other values of $\alpha$, the class of tests generated by $\mathcal{F}$ will not contain a similar test, though nonrandom similar tests may exist outside of this class. Tests at other values of $\alpha$ are typically of interest for two reasons: using multiple testing adjustment procedures and defining $p$-values. We discuss the latter in the Supplementary Materials. The restriction that $\alpha$ must be a unit fraction is not a limitation for the Bonferroni correction procedure, which involves dividing the familywise error rate $\alpha$ by the number of tests, provided the familywise error rate $\alpha$ itself is a unit fraction. However, it is a limitation for the Benjamini–Hochberg procedure, which involves multiplying the false discovery rate $\alpha$ by a sequence of rational numbers, only some of which are unit fractions. We discuss the latter procedure in Section 5.
We now define an extension of the minimax optimal test to non-unit fraction values of $\alpha\in(0,1)$. When $\alpha$ is not a unit fraction, then this extension will not yield a minimax optimal test or a similar test, as similar tests do not exist in this class of tests for such values. However, they do preserve type 1 error, and have type 1 error that approximates $\alpha$ under the least-favorable distribution in the null hypothesis very closely as $\alpha$ goes to zero. Let $b_{-1}:=0$ and $b_{k}:=\Phi^{-1}\{1-(\lfloor\alpha^{-1}\rfloor-k)\alpha/2\}$ for $k=0,\ldots,\lfloor\alpha^{-1}\rfloor$ so that $0=b_{-1}\leq b_{0}<b_{1}<\ldots<b_{\lfloor\alpha^{-1}\rfloor}=+\infty$. We define the rejection region for any $\alpha\in(0,1)$ to be $R_{mm}^{\prime}:=\{\bigcup_{k=0}^{\lfloor\alpha^{-1}\rfloor}(b_{k-1},b_{k})\times(b_{k-1},b_{k})\}\cup\{\bigcup_{k=0}^{\lfloor\alpha^{-1}\rfloor}(b_{k-1},b_{k})\times(-b_{k},-b_{k-1})\}\cup\{\bigcup_{k=0}^{\lfloor\alpha^{-1}\rfloor}(-b_{k},-b_{k-1})\times(b_{k-1},b_{k})\}\cup\{\bigcup_{k=0}^{\lfloor\alpha^{-1}\rfloor}(-b_{k},-b_{k-1})\times(-b_{k},-b_{k-1})\}$. When $\alpha$ is a unit fraction, $R^{\prime}_{mm}$ coincides with $R_{mm}$, hence $R^{\prime}_{mm}$ generates the minimax optimal test discussed earlier. Otherwise, the test generated by $R^{\prime}_{mm}$ corresponds closely to the minimax optimal test in that if $\delta_{y}^{*}=0$, then given $Z^{x}_{*}=z$ with $z\notin[-b_{0},b_{0}]$, the probability of being in the rejection region is exactly $\alpha$. Precisely:
@@
$\lfloor\alpha^{-1}\rfloor\alpha^{2}+(1-\lfloor\alpha^{-1}\rfloor\alpha)^{2}$ $=\inf_{\bm{\delta}^{*}:\delta^{*}_{x}\delta^{*}_{y}=0}\mathrm{Pr}_{\bm{\delta}^{*}}\{(Z_{*}^{x},Z_{*}^{y})\in R_{mm}^{\prime}\}$ (7)
$\leq\sup_{\bm{\delta}^{*}:\delta^{*}_{x}\delta^{*}_{y}=0}\mathrm{Pr}_{\bm{\delta}^{*}}\{(Z_{*}^{x},Z_{*}^{y})\in R_{mm}^{\prime}\}=\alpha.$
@@ 3.3 Asymptotic approximation considerations
The results we have given so far have been in terms of the limit distribution, i.e., that of $(Z^{x}_{*},Z^{y}_{*})$. However, if the joint convergence of $(\hat{\delta}_{x},\hat{\delta}_{y})$ is uniform in $\bm{\delta}$ as in (1), then the power $\beta_{n}(\delta_{n}):=\mathrm{Pr}_{\hat{\delta}_{x},\hat{\delta}_{y}}\{(Z^{x}_{n},Z^{y}_{n})\in R_{mm}\}$ for the observed data distribution will converge uniformly to $\beta_{*}(\delta^{*}):=\mathrm{Pr}_{\bm{\delta}^{*}}\{(Z^{x}_{*},Z^{y}_{*})\in R_{mm}\}$, i.e., that of the limiting distribution. Thus, for any arbitrarily small tolerance $\epsilon>0$, there is an $N$ such that $\beta_{n}(\delta_{n})$ will be within $\epsilon$ of $\beta_{*}(\delta^{*})$ for all $\bm{\delta}$ and $n\geq N$. In particular, this means violations of type 1 error can be made arbitrarily small uniformly over $H_{0}$ given a sufficiently large sample size. The asymptotic approximation can be improved for a given finite sample $n$ by replacing all instances of the normal distribution with a $t$ distribution with $(n-1)$ degrees of freedom.